# On the self-consistency of compact objects in Lorentz-violating gravity theories

### v1

@@ Abstract
Antisymmetric rank-2 tensor constraints. Now, let us consider the Einstein-antisymmetric rank-2 tensor model ($\xi_{2}\neq 0$ and $\xi_{3}\neq 0$) and that the dynamics of the antisymmetric rank-2 tensor field is frozen in its VEV sate, $\langle B_{\mu\nu}\rangle=b_{\mu\nu}$. As demonstrated in Ref. [9], the VEV $b_{\mu\nu}$ can be decomposed as $b_{\mu\nu}=\tilde{E}_{[\mu}v_{\nu]}+\epsilon_{\mu\nu\alpha\beta}v^{\alpha}\tilde{B}^{\beta}$, where the background vectors $\tilde{E}_{\mu}$ and $\tilde{B}_{\mu}$ can be interpreted as pseudo-electric and pseudo-magnetic fields, respectively, and $v^{\mu}$ is a timelike 4-vector. The pseudo-fields $\tilde{E}_{\mu}$ and $\tilde{B}_{\mu}$ are spacelike, i.e., $\tilde{E}_{\mu}v^{\mu}=\tilde{B}_{\mu}v^{\mu}=0$.
@@
$4\langle V_{X}\rangle b^{\mu\nu}+\frac{2\xi_{2}}{\kappa}b_{\alpha}{}^{[\mu}R^{\nu]\alpha}-\frac{2\xi_{3}}{\kappa}b_{\alpha\beta}R^{\alpha\beta\mu\nu}=0,$ (19)
@@
$\xi_{2}\left(r\Omega A^{\prime\prime}-rA^{\prime}\Omega^{\prime}+2\Omega A^{\prime}-2A\Omega^{\prime}\right)$ (20)
$+\xi_{3}\left(2r\Omega A^{\prime\prime}-2rA^{\prime}\Omega^{\prime}\right)+4\kappa r\langle V_{X}\rangle\Omega^{3}=0.$
@@
Particularly, if $\Omega^{2}(r)$ is a constant, for instance $\Omega^{2}(r)=\omega^{2}$, the constraint (20) reduces to
$rA^{\prime\prime}(\xi_{2}+2\xi_{3})+2\xi_{2}A^{\prime}+4\kappa r\langle V_{X}\rangle\omega^{2}=0.$ (21)
By considering that the potential that drives the breaking of the Lorentz symmetry extremizes at the VEV state, for example a quadratic one, both $\langle{V}\rangle$ and $\langle{V}_{X}\rangle$ vanish. Thus, the above equation becomes $rA^{\prime\prime}(\xi_{2}+2\xi_{3})+2\xi_{2}A^{\prime}=0$. Consequently, to satisfy the constraint imposed by $b_{\mu\nu}$, the redshift function must conform to
@@
where $\gamma=1-{2\xi_{2}}/(\xi_{2}+2\xi_{3})$ and ${c}_{1,2}$ are integration constants that cannot be fixed solely from the constraint (8). It follows that depending on the combination of the coupling constants, different families of functions $A(r)$ are allowed by the model. For instance, if $\xi_{3}=0$, $\gamma=-1$, and consequently $A(r)=c_{1}-c_{2}/r$, which has a Schwarzschild-like form. On the other hand, if $\xi_{2}=0$, $\gamma=1$, and consequently $A(r)=c_{1}+c_{2}r$, where the linear term can be understood as a sort of Rindler-type acceleration [19, 20]. Since $\gamma$ can take arbitrary real values, even an (A)dS-like redshift function is allowed, for instance if $\xi_{3}=-3\xi_{2}/2$, $\gamma=2$, and $A=c_{1}+c_{2}r^{2}/2$. Although redshift functions with arbitrary exponents can indeed satisfy the above constraint, finding reasonable matter configurations that give support to such configurations is difficult, therefore they have limited physical relevance.
We emphasize that black hole solutions in the Einstein-antisymmetric rank-2 tensor system—under the VEV state considered here—have already been reported in the literature. Specifically, the black hole in Ref. [11], defined by the metric functions $A(r)={1}/{(1-l_{T})}-{2M}/{r}$ and $B(r)=A(r)$, and the solution in Ref. [12], described by $A(r)=1-{2M}/{r}$ and $B(r)={A(r)}/{(1-l_{T})}$, both conform to the constraint (21). Here, $l_{T}=b^{2}\tilde{\xi}_{2}/2$ encodes Lorentz-violating effects, while $M$ represents the mass of the compact object. Therefore, these black holes are self-consistent within the Einstein-antisymmetric rank-2 tensor model. Conversely, in Ref. [17], it was reported a vacuum black hole solution, described by $A(r)=1-2M/r+\Upsilon/r^{2/l_{T}}$ and $B(r)=A(r)$, with $\Upsilon$ being a constant, that, as we can see from Eq. (22), does not conform to the constraint (21). Hence, this black hole is not self-consistent within the considered model and therefore cannot be cast as a solution of it.
@@
$A(r)=\tilde{c}_{1}+\frac{\tilde{c}_{2}r^{\gamma}}{\gamma}-\frac{\omega^{2}\tilde{\Lambda}_{e}}{3}r^{2}$ (23)
in order to respect the anisotropies induced by $b_{\mu\nu}$. Here $\tilde{c}_{1,2}$ are integration constants that cannot be fixed solely from the constraint (8) and $\tilde{\Lambda}_{e}=\frac{\kappa\lambda}{3\xi_{2}+{2}\xi_{3}/{3}}$, with $\xi_{3}\neq-3\xi_{2}/2$. Particularly interesting cases cases are given by $\xi_{3}=0$, that leads to $\gamma=-1$, $\tilde{\Lambda}_{e}=\kappa\lambda/2\xi_{2}$ and a a Kottler-like redshift function, and $\xi_{2}=0$, that leads to $\gamma=1$, $\tilde{\Lambda}_{e}=\kappa\lambda/3\xi_{3}$ and a redshift function with an effective cosmological constant and a Rindler-type acceleration, respectively. We point out that—under the VEV state considered here—static and spherically symmetric vacuum solutions of the Einstein-antisymmetric rank-2 tensor system with cosmological constant have been reported in the literature [11, 12]. We remark that these solutions satisfy the constraint (23).
@@
$\xi_{3}\{2r[s(r)+rs^{\prime}(r)]\Phi^{\prime}(r)+4r^{2}[r-s^{\prime}(r)][\Phi^{\prime}(r)^{2}-\Phi^{\prime\prime}(r)]\}$
$+{\xi}_{2}\{s(r)-rs^{\prime}(r)-[s(r)+r(s^{\prime}(r)-2)]r\Phi^{\prime}(r)\}$
$0\displaystyle 2{\xi}_{2}r^{2}[r-s(r)][\Phi^{\prime}(r)^{2}+\Phi^{\prime\prime}(r)]+4r^{3}\kappa\langle{V}_{X}\rangle=$ 0 . (24)
@@
In particular, let us probe if tideless wormholes are self-consistent objects in such a system. By considering $\Phi(r)=0$, the above constraint simplifies to $\xi_{2}[s(r)-rs^{\prime}(r)]=-4r^{3}\kappa\langle V_{X}\rangle$. Therefore, considering that the potential extremizes at the VEV state, one finds that tideless wormholes are not self-consistent objects within the Einstein-antisymmetric rank-2 tensor model, once the areal radius $r(x)$ cannot have a local minimum, similarly as the constraint present in the Einstein-bumblebee model. However, by considering a linear potential, $V(X)=\frac{1}{2}\lambda X$, one is able to find self-consistent tideless wormholes. Specifically, their shape function must behave as $s(r)=(1-a^{2}\tilde{\Lambda}_{e})r+\tilde{\Lambda}_{e}r^{3}$. In particular, our results show that the tideless wormhole reported in Ref. [18], although it satisfies the corresponding Einstein equations, is not self-consistent within the Lorentz-violating curvature-coupled antisymmetric rank-2 tensor model and cannot be cast as a solution of this model.
@@
Remarkably, the whole term multiplying $\xi_{3}$ in Eq. (On the self-consistency of compact objects in Lorentz-violating gravity theories) vanishes for tideless configurations, regardless of the form of the shape function. It implies that, in Riemann-coupled antisymmetric rank-2 tensor models ($\xi_{2}=0$ and $\xi_{3}\neq 0$), considering self-interaction potential that extremizes at the VEV state, the constraint (On the self-consistency of compact objects in Lorentz-violating gravity theories) is identically satisfied. However, if instead a linear self-interaction potential is considered, tideless wormholes are ruled out by the above constraint. Notably, a tideless wormhole of this system was reported in Ref. [15]. The authors find a tideless wormhole configuration sourced by an ideal fluid, considering a potential that extremizes at the VEV state, which by the above analysis is a self-consistent object of the Riemann-coupled antisymmetric rank-2 tensor model.
@@
We remark that the solution reported in Ref. [15] is not the unique possible tideless wormhole supported within this framework, and other tideless wormholes, associated to different shape functions, can be found. Surprisingly, a Lorentz-violating version of the (massless) Ellis-Bronnikov wormhole, namely a tideless wormhole with shape function $s(r)\propto a^{2}/r$, is still missing in the literature of Lorentz-violating bumbeblee or antisymmetric rank-2 tensors fields, and due to our analysis, such wormhole is self consistent within the latter Lorentz-violating gravity framework. A derivation of such wormhole and other viable wormholes that satisfy the above discussed criteria is currently underway.
@@
Final words. The self-consistency of compact objects in Lorentz-violating gravity theories demands not only the fulfillment of Einstein and matter field equations but also the satisfaction of the Lorentz-violating field equations. These equations can impose stringent geometric constraints in vacuum states, which we argue may rule out the existence of entire families of compact object solutions in such frameworks. By analyzing two prominent models—spontaneous Lorentz symmetry breaking via a bumblebee vector field and an antisymmetric rank-2 tensor field—we demonstrate that common assumptions, such as freezing Lorentz-violating fields at their VEVs while vanishing their field strength, may obstruct rather than enable the emergence of certain compact objects. Our results challenge the physical admissibility of several previously reported solutions, including black holes and wormholes, as their failure to satisfy the Lorentz-violating field equations leads to inconsistencies such as non-conservation of the total energy-momentum tensor.
@@
A promising avenue to circumvent these constraints lies in coupling Lorentz-violating fields to matter. Such couplings can modify the dynamics of these fields in their vacuum state, potentially permitting a broader class of solutions through induced anisotropies. For instance, coupling a $U(1)$ gauge field to Lorentz-violating fields has successfully generated self-consistent Lorentz-violating analogues of Reissner-Nordström black holes [21, 22]. This approach opens a pathway to extend well-known GR solutions into Lorentz-violating frameworks while preserving self-consistency. In particular, Lorentz-violating analogues of the Ellis-Bronnikov wormholes, excluded in the Einstein-bumblebee model considered here under vacuum conditions, could become viable with appropriate matter couplings. A detailed investigation of such wormhole solutions in the Einstein-bumblebee model is in progress, with results to be reported in due course.
@@ Acknowledgements.
The authors would like to acknowledge Fundação de Amparo à Pesquisa e ao Desenvolvimento Científico e Tecnológico do Maranhão (FAPEMA), Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES) – Finance Code 001, from Brazil, for partial financial support. L.A.L is supported by FAPEMA BPD- 08975/24.

### v2

@@ V Antisymmetric rank-2 tensor constraints
Now, let us consider the Einstein-antisymmetric rank-2 tensor model ($\tilde{\xi}_{1,2,3}=0$) and that the dynamics of the antisymmetric rank-2 tensor field is frozen in its VEV state, $\langle B_{\mu\nu}\rangle=b_{\mu\nu}$. As demonstrated in Ref. [45], the VEV $b_{\mu\nu}$ can be decomposed as $b_{\mu\nu}=\tilde{E}_{[\mu}v_{\nu]}+\epsilon_{\mu\nu\alpha\beta}v^{\alpha}\tilde{B}^{\beta}$, where the background vectors $\tilde{E}_{\mu}$ and $\tilde{B}_{\mu}$ can be interpreted as pseudo-electric and pseudo-magnetic fields, respectively, and $v^{\mu}$ is a timelike 4-vector. The pseudo-fields $\tilde{E}_{\mu}$ and $\tilde{B}_{\mu}$ are spacelike, i.e., $\tilde{E}_{\mu}v^{\mu}=\tilde{B}_{\mu}v^{\mu}=0$.
@@
$4\langle V_{X}\rangle b^{\mu\nu}-\frac{2\xi_{1}}{\kappa}Rb^{\mu\nu}+\frac{2\xi_{2}}{\kappa}b_{\alpha}{}^{[\mu}R^{\nu]\alpha}-\frac{2\xi_{3}}{\kappa}b_{\alpha\beta}R^{\alpha\beta\mu\nu}=0,$ (28)
@@
$\frac{\xi_{1}}{r}[-2r^{2}A^{\prime}\Omega^{\prime}+2r\Omega\left(rA^{\prime\prime}+4A^{\prime}\right)+4A\left(\Omega-2r\Omega^{\prime}\right)-4\Omega^{3}]$
$\xi_{2}[\left(rA^{\prime}+2A\right)\Omega^{\prime}-\Omega\left(rA^{\prime\prime}+2A^{\prime}\right)]$ (29)
$+2\xi_{3}r\left(\Omega A^{\prime\prime}-A^{\prime}\Omega^{\prime}\right)+4\kappa r\langle V_{X}\rangle\Omega^{3}=0.$
@@
Particularly, if $\Omega^{2}(r)$ is a constant, for instance $\Omega^{2}(r)=\omega^{2}$, the constraint (V) reduces to
$\frac{2\xi_{1}\left(r^{2}A^{\prime\prime}+4rA^{\prime}+2A+2\omega^{2}\right)}{r}+rA^{\prime\prime}(\xi_{2}+2\xi_{3})$
$+2\xi_{2}A^{\prime}+4\kappa r\langle V_{X}\rangle\omega^{2}=0.$ (30)
We point out that the analytical investigation of the above constraint is challenging if $\xi_{1,2,3}$ are simultaneously non-zero. Thus, for simplicity, let us first address the vanishing ${\xi}_{1}$ case. By considering that the potential that drives the breaking of the Lorentz symmetry extremizes at the VEV state, for example a quadratic one, both $\langle{V}\rangle$ and $\langle{V}_{X}\rangle$ vanish. Thus, the above equation becomes $rA^{\prime\prime}(\xi_{2}+2\xi_{3})+2\xi_{2}A^{\prime}=0$. Consequently, to satisfy the constraint imposed by $b_{\mu\nu}$, the redshift function must conform to
@@
where $\gamma=1-{2\xi_{2}}/(\xi_{2}+2\xi_{3})$ and ${c}_{1,2}$ are integration constants that cannot be fixed solely from the constraint (II). It follows that depending on the combination of the coupling constants, different families of functions $A(r)$ are allowed by the model. For instance, if $\xi_{3}=0$, $\gamma=-1$, and consequently $A(r)=c_{1}-c_{2}/r$, which has a Schwarzschild-like form. On the other hand, if $\xi_{2}=0$, $\gamma=1$, and consequently $A(r)=c_{1}+c_{2}r$, where the linear term can be understood as a sort of Rindler-type acceleration [58, 59]. Since $\gamma$ can take arbitrary real values, even an (A)dS-like redshift function is allowed, for instance if $\xi_{3}=-3\xi_{2}/2$, $\gamma=2$, and $A=c_{1}+c_{2}r^{2}/2$. Although redshift functions with arbitrary exponents can indeed satisfy the above constraint, finding reasonable matter configurations that give support to such configurations is difficult, therefore they have limited physical relevance.
We emphasize that black hole solutions in the Einstein-antisymmetric rank-2 tensor system—under the VEV state considered here—have already been reported in the literature. Specifically, the black hole in Ref. [31], defined by the metric functions $A(r)={1}/{(1-l_{T})}-{2M}/{r}$ and $B(r)=A(r)$, and the solution in Ref. [32], described by A(r)=12M/rA(r)=1-{2M}/{r} and $B(r)={A(r)}/{(1-l_{T})}$, both conform to the constraint (V). Here, $l_{T}=b^{2}\tilde{\xi}_{2}/2$ encodes Lorentz-violating effects, while $M$ represents the mass of the compact object. Therefore, these black holes are self-consistent within the Einstein-antisymmetric rank-2 tensor model. Remarkably, these solutions share an identical functional form with vacuum-bumblebee black holes [30], differing only by a sign change in the Lorentz-violating parameter. Consequently, the same geometric consequences apply: spontaneous local Lorentz and diffeomorphism symmetry breaking manifests as a conical structure with deficit angle $\Delta=\pi^{2}/\sqrt{1-l_{T}}$, resulting in non-flat asymptotics. The conical geometry in the equatorial plane similarly modifies weak-field phenomena like light deflection [31].
Conversely, in Ref. [52], it was reported a vacuum black hole solution, described by $A(r)=1-2M/r+\Upsilon/r^{2/l_{T}}$ and $B(r)=A(r)$, with $\Upsilon$ being a constant, that, as we can see from Eq. (31), does not conform to the constraint (V). Hence, this black hole is not self-consistent within the considered model and therefore cannot be cast as a solution of it.
@@
$A(r)=\tilde{c}_{1}+\frac{\tilde{c}_{2}r^{\gamma}}{\gamma}-\frac{\omega^{2}\Lambda_{e}}{3}r^{2}$ (32)
in order to respect the preferred spacetime direction induced by $b_{\mu\nu}$. Here $\tilde{c}_{1,2}$ are integration constants that cannot be fixed solely from the constraint (II) and $\Lambda_{e}=\frac{\kappa\lambda}{3\xi_{2}+{2}\xi_{3}/{3}}$, with $\xi_{3}\neq-3\xi_{2}/2$. Particularly interesting cases are given by $\xi_{3}=0$, that leads to $\gamma=-1$, $\Lambda_{e}=\kappa\lambda/2\xi_{2}$ and a Kottler-like redshift function, and $\xi_{2}=0$, that leads to $\gamma=1$, $\Lambda_{e}=\kappa\lambda/3\xi_{3}$ and a redshift function with an effective cosmological constant and a Rindler-type acceleration, respectively. We point out that—under the VEV state considered here—static and spherically symmetric vacuum solutions of the Einstein-antisymmetric rank-2 tensor system with cosmological constant have been reported in the literature [31, 32]. We remark that these solutions satisfy the constraint (28).
For completeness, let us also examine the Einstein-antisymmetric rank-2 tensor model with exclusively non-minimal Ricci scalar coupling ($\xi_{1}\neq 0$). In this case, the constraint (V) reduces to
$\frac{\xi_{1}\left(r^{2}A^{\prime\prime}+4rA^{\prime}+2A+2\omega^{2}\right)}{r}=-2\kappa r\langle V_{X}\rangle\omega^{2}.$ (33)
We note that the above equation shares the same form as Eq. (20), implying that the lapse functions derived for the bumblebee case also satisfy this constraint. Similar to the bumblebee scenario, no compact object solutions with these lapse function structures have been reported as solutions of this model, suggesting that this can be a novel class of solutions worthy of detailed study.
@@
$2\xi_{1}r\{-s^{\prime}\left(r\Phi^{\prime}+2\right)+(4r-3s)\Phi^{\prime}+2r(r-s)\left(\Phi^{\prime\prime}+\Phi^{\prime 2}\right)\}$
$+\xi_{3}\{2r[s+rs^{\prime}]\Phi^{\prime}+4r^{2}[r-s^{\prime}][\Phi^{\prime 2}-\Phi^{\prime\prime}]\}$
$+{\xi}_{2}\{s-rs^{\prime}-[s+r(s^{\prime}-2)]r\Phi^{\prime}\}$
$+2{\xi}_{2}r^{2}[r-s][\Phi^{\prime 2}+\Phi^{\prime\prime}]+4r^{3}\kappa\langle{V}_{X}\rangle=0.$ (34)
@@
In particular, let us probe whether tideless wormholes are self-consistent objects in such a system. Considering $\Phi=0$, the above constraint simplifies to $4\xi_{1}rs^{\prime}+\xi_{2}[s-rs^{\prime}]=-4r^{3}\kappa\langle V_{X}\rangle$. Therefore, considering a self-interaction potential that extremizes at the VEV state, this constraint takes a form similar to that of the bumblebee model, and one can check that the same conclusion applies. Furthermore, one can also find self-consistent tideless wormhole solutions within this model by considering a linear potential. Remarkably, their structure closely resembles that of the solutions permitted in the bumblebee case. In particular, our results show that the tideless wormhole reported in Ref. [53], although it satisfies the corresponding Einstein equations, is not self-consistent within the considered Lorentz-violating curvature-coupled antisymmetric rank-2 tensor model and cannot be cast as a solution of this model.
@@
Remarkably, the whole term multiplying $\xi_{3}$ in Eq. (V) vanishes for tideless configurations, regardless of the form of the shape function. It implies that, in Riemann-coupled antisymmetric rank-2 tensor models ($\xi_{2}=0$ and $\xi_{3}\neq 0$), considering self-interaction potential that extremizes at the VEV state, the constraint (V) is identically satisfied. However, if instead a linear self-interaction potential is considered, tideless wormholes are ruled out by the above constraint. Notably, a tideless wormhole of this system was reported in Ref. [35]. The authors find a tideless wormhole configuration sourced by an ideal fluid, considering a potential that extremizes at the VEV state, which by the above analysis is a self-consistent object of the Riemann-coupled antisymmetric rank-2 tensor model.
@@
We remark that the solution reported in Ref. [35] is not the unique possible tideless wormhole supported within this framework, and other tideless wormholes, associated to different shape functions, can be found. Surprisingly, a Lorentz-violating version of the (massless) Ellis-Bronnikov wormhole, namely a tideless wormhole with shape function $s(r)\propto a^{2}/r$ [60, 61], is still missing in the literature of Lorentz-violating bumbeblee or antisymmetric rank-2 tensors fields, and due to our analysis, such wormhole is self consistent within the latter Lorentz-violating gravity framework. A derivation of such wormhole and other viable wormholes that satisfy the above discussed criteria is currently underway.
VI Final words
@@
The self-consistency of compact objects in theories with spontaneous local Lorentz and diffeomorphism violation demands not only the fulfillment of Einstein and matter field equations but also the satisfaction of the Lorentz-violating field equations. These equations can impose stringent geometric constraints in vacuum states, which we argue may rule out the existence of entire families of compact object solutions in such frameworks. By analyzing two prominent models— spontaneous local Lorentz and diffeomorphism symmetry breaking via a bumblebee vector field and an antisymmetric rank-2 tensor field—we demonstrate that common assumptions, such as freezing Lorentz-violating fields at their VEVs while vanishing their field strength, may obstruct rather than enable the emergence of certain compact objects. Our results challenge the physical admissibility of several previously reported solutions, including black holes and wormholes, as their failure to satisfy the Lorentz-violating field equations leads to inconsistencies such as non-conservation of the total energy-momentum tensor. We emphasize that by avoiding specific assumptions about the VEVs and allowing the background field dynamics to evolve naturally, enables a broader class of compact object solutions [34].
@@
A promising avenue to circumvent these constraints lies in coupling Lorentz-violating fields to matter. Such couplings can modify the dynamics of these fields in their vacuum state, potentially permitting a broader class of solutions. For instance, coupling a $U(1)$ gauge field to Lorentz-violating fields has successfully generated self-consistent Lorentz-violating analogues of Reissner-Nordström black holes [62, 63]. This approach opens a pathway to extend well-known GR solutions into Lorentz-violating frameworks while preserving self-consistency. In particular, Lorentz-violating analogues of the Ellis-Bronnikov wormholes could become viable with appropriate matter couplings. A detailed investigation of such wormhole solutions in the Einstein-bumblebee and in the Einstein-antisymmetric rank-2 tensor models is in progress, with results to be reported in due course.
@@ Acknowledgements.
The authors would like to acknowledge Fundação de Amparo à Pesquisa e ao Desenvolvimento Científico e Tecnológico do Maranhão (FAPEMA), Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES) – Finance Code 001, from Brazil, for partial financial support. R.B.M. is supported by CNPq/PDJ 151250/2024-3. L.A.L is supported by FAPEMA BPD- 08975/24.